ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 126, 2001

Adams, T. M. and A. B. Nobel Finitary reconstruction of a measure preserving transformation

309

Ban, D. and C. Jantzen The Langlands classification for nonconnected p-adic groups

239

Bump, D., S. Friedberg and D. Ginzburg On the cubic Shimura lift for PGL_{3}

289

Cilleruelo, J. and C. Trujillo Infinite B_{2}[g] sequences  263
Dani, S. G. On ergodic ℤ^{d}-actions on Lie groups by automorphisms  327

Dasgupta, S., G. Károlyi, O. Serra and B. Szegedy Transversals of additive Latin squares

17

Friedberg, S. See D. Bump

Gelbart, S., H. Jacquet and J. Rogawski Generic representations for the unitary group in three variables

173

Gelbord, B. and R. Meshulam Spaces of p-vectors of bounded rank 129

Ginzburg, D. See D. Bump

Guralnick, R. and P. Shumyatsky Derived subgroups of fixed points

345

Jacquet, H. See S. Gelbart

Jantzen, C. See D. Ban

Károlyi, G. See S. Dasgupta

Larsen, M. How often is 84(g − 1) achieved?  1

Meshulam, R. See B. Gelbord

Niţică, V. A note about topologically transitive cylindrical cascades 141

Nobel, A. B. See T. M. Adams

Rogawski, J. See S. Gelbart

du Sautoy, M. A nilpotent group and its elliptic curve: Nonuniformity of local zeta functions of groups

269

Serra, O. See S. Dasgupta

Shelah, S. Categoricity of an abstract elementary class in two successive cardinals

29

Shields, P. C. Loose almost blowing-up and related properties  157

Shumyatsky, P. See R. Guralnick Szegedy, B. See S. Dasgupta Trujillo, C. See J. Cilleruelo

Wulbert, D. A general Hobby–Rice Theorem and cake cutting  363